The horizontal distance between a point of equilibrium on a wave and the  middle of a crest is 2.50 m. What is the period if the wave travels 20.0 m  in 35.0 s?


1
Expert's answer
2021-06-17T15:27:40-0400

Gives

A=2.50m

x=2.50m

t=35.0sec

"v=Aw"

"v=2.50\times\frac{2\pi}{T}\rightarrow equation(1)"

"v=\frac{x}{T}=\frac{20}{35}\rightarrow(2)"

Equation (1) and(2) are equal

"\frac{20}{35}=2.50\times\frac{2 \pi}{T}"

T=28 sec